5. Trapezoid ABCD has diagonals AC and BD. If mZABD = 5x-10 and mZCDB = 3x+2, what is the value of x?
thanks
Answer #1

In the trapezium ABCD, AB is parallel to BD. Therefore mABD = mCDB (alternate angles between 2 parallel lines are equal).

Therefore 5x - 10 = 3x + 2

Solving, we get 5x - 3x = 10 + 2

2x = 12

Therefore x = 6